# The Label Horizon Paradox: Rethinking Supervision Targets in Financial Forecasting

### v2

@@ 3.2 Theoretical Framework
The expected performance is determined by the net balance of two competing accumulation processes (ignoring a constant term related to target variance):
@@
1. Information Gain reflects the growth of valid signal in the label. It increases as $\delta$ extends, but naturally saturates as $\alpha(\delta)\to 1$ (when information is fully priced).
@@ 3.3 Generalization to Deep Learning
To empirically validate this theoretical connection, we specifically examine Scenario 1. Given the interday nature of this task, the market has ample time to absorb the prior day’s information, causing the signal embedded in the input features to be priced in immediately upon the market open. Consequently, the signal realization saturates almost immediately ($\alpha(\delta)\approx 1$). This leads to the following approximation:
@@ 3.4 Mechanism Analysis
Scenario 1 (Daily): Interday signals typically saturate early (e.g., at the market open). Extending $\delta$ beyond this point yields vanishing information gain ($\alpha^{\prime}\approx 0$) while the noise penalty continues to grow linearly. Thus, the gradient is strictly negative, driving the optimal horizon to be much smaller than the target ($\delta^{*}\ll\Delta$).
@@ A.5.1 Predictive Model Training
A.5.2 Bi-level Optimization Framework.
@@ C.1.1 The Standard Static APT
$r^{\Delta}=\mathbf{w}^{*\top}\mathbf{s}+\epsilon,$ (16)
where $\mathbf{s}$ represents factor exposures, $\mathbf{w}^{*}$ represents factor risk premia, and $\epsilon$ is the idiosyncratic noise. This model is typically static—it describes the return over a single, undefined ”period” where information is assumed to be fully reflected.
@@ C.1.2 Temporal Extension
$r^{\delta}=\underbrace{\alpha(\delta)}_{\text{Signal Realization}}\mathbf{w}^{*\top}\mathbf{s}+\underbrace{\epsilon^{\delta}}_{\text{Accumulated Noise}}$ (17)
@@
    Signal Realization Process ($\alpha(\delta)$): Unlike the standard APT, which assumes equilibrium (i.e., information is fully priced), we acknowledge that information incorporation takes time. Let $\alpha:[1,\Delta]\to(0,1]$ be a monotonically increasing function representing the degree of price discovery.
    • $\alpha(\delta)<1$ implies partial underreaction or gradual diffusion of information.
    • $\alpha(\delta)=1$ implies the signal is fully realized.
@@
    Here, $\sigma^{2}$ is the accumulation rate of idiosyncratic volatility, and $\delta_{0}>0$ represents intrinsic microstructure noise that exists even at $\delta\to 1$. While this assumption is generally reasonable at short horizons such as intraday (minute-level) and daily intervals, it may become less accurate over much longer horizons where structural breaks and slow-moving factors dominate. Since our study focuses on short-term stock forecasting at minute and daily frequencies, the random-walk-based noise model is well aligned with the forecasting regimes of interest.
@@ C.2 The Learning Setup: Finite-Sample OLS
$\hat{\mathbf{w}}_{\delta}=(\mathbf{S}^{\top}\mathbf{S})^{-1}\mathbf{S}^{\top}\mathbf{r}^{\delta}=\alpha(\delta)\mathbf{w}^{*}+(\mathbf{S}^{\top}\mathbf{S})^{-1}\mathbf{S}^{\top}\boldsymbol{\epsilon}^{\delta}$ (19)
@@
$\hat{\mathbf{w}}_{\delta}=\alpha(\delta)\mathbf{w}^{*}+\mathbf{z}_{\delta},\quad\text{where }\mathbf{z}_{\delta}\sim\mathcal{N}\left(\mathbf{0},\frac{\sigma^{2}(\delta+\delta_{0})}{N}\mathbf{I}_{d}\right)$ (20)
@@ C.3 Derivation of Generalization Performance (Final IC)
    $r^{\Delta}=\mathbf{w}^{*\top}\mathbf{s}+\epsilon^{\Delta}.$
@@
    $\text{Cov}(\hat{y}^{\delta},r^{\Delta})$ $=\text{Cov}\big(\alpha(\delta)\mathbf{w}^{*\top}\mathbf{s}+\mathbf{z}_{\delta}^{\top}\mathbf{s},\,\mathbf{w}^{*\top}\mathbf{s}+\epsilon^{\Delta}\big)$ (21)
    $=\alpha(\delta)\,\text{Var}(\mathbf{w}^{*\top}\mathbf{s})+\underbrace{\text{Cov}(\mathbf{z}_{\delta}^{\top}\mathbf{s},\mathbf{w}^{*\top}\mathbf{s})}_{=0}+\underbrace{\text{Cov}(\alpha(\delta)\mathbf{w}^{*\top}\mathbf{s},\epsilon^{\Delta})}_{=0}+\underbrace{\text{Cov}(\mathbf{z}_{\delta}^{\top}\mathbf{s},\epsilon^{\Delta})}_{=0}$ (22)
    $=\alpha(\delta)\cdot 1=\alpha(\delta).$ (23)
@@
    $\text{Cov}(\hat{y}^{\delta},r^{\Delta})=\text{Cov}(\alpha(\delta)\mathbf{w}^{*\top}\mathbf{s},\mathbf{w}^{*\top}\mathbf{s})=\alpha(\delta).$ (24)
@@
    $V_{\text{target}}=\text{Var}(r^{\Delta})=1+\sigma^{2}(\Delta+\delta_{0}).$ (26)
@@
$J(\delta)\triangleq\text{IC}_{\text{final}}^{2}(\delta)=\frac{\text{Cov}(\hat{y}^{\delta},r^{\Delta})^{2}}{V_{\text{est}}(\delta)V_{\text{target}}}=\frac{\alpha(\delta)^{2}}{\left[\alpha(\delta)^{2}+K(\delta+\delta_{0})\right]\cdot V_{\text{target}}},$ (27)
@@ C.4 Proof of the Paradox
To determine the optimal training horizon $\delta^{*}$, we analyze the behavior of the expected squared IC, $J(\delta)$. Since the target variance $V_{\text{target}}$ is a constant scalar independent of the training horizon $\delta$, maximizing $J(\delta)$ is equivalent to maximizing the log-likelihood of $\alpha(\delta)^{2}/\left[\alpha(\delta)^{2}+K(\delta+\delta_{0})\right]$.
@@ C.5 Detailed Mechanism Analysis of Different Horizon Scenarios
Scenario 3: Intraday 90-minute Prediction (Hump-Shaped). When the target window extends to 90 minutes, we observe the characteristic ”Label Horizon Paradox.” This scenario represents the transition between active information diffusion and signal saturation.
@@ C.6 Decomposition under Interday Prediction Scenario
Mathematically, this corresponds to the regime of rapid price discovery. The signal realization function satisfies $\alpha(\delta)\approx 1$ and $\alpha^{\prime}(\delta)\approx 0$ for any horizon $\delta$ extending beyond the market open.
Under this condition (where $\alpha(\delta)=1$), and recalling that the idiosyncratic noise follows a nested random-walk structure such that $\text{Cov}(\epsilon^{\delta},\epsilon^{\Delta})=\text{Var}(\epsilon^{\delta})$, the covariance between the proxy label $r^{\delta}$ and the target $r^{\Delta}$ simplifies to the variance of the proxy itself:
$\text{Cov}(r^{\delta},r^{\Delta})=\text{Var}(\mathbf{w}^{*\top}\mathbf{s})+\text{Cov}(\epsilon^{\delta},\epsilon^{\Delta})=1+\sigma^{2}(\delta+\delta_{0})\equiv V_{\text{proxy}}(\delta).$ (36)
This equality is crucial, as it allows for an exact multiplicative decomposition of the final performance. The final Information Coefficient (IC) can be factorized into the product of the model’s fit to the proxy and the proxy’s correlation with the target:
$\underbrace{\frac{1}{\sqrt{V_{\text{est}}}\sqrt{V_{\text{proxy}}}}}_{\rho(\hat{y}^{\delta},r^{\delta})}\times\underbrace{\frac{V_{\text{proxy}}}{\sqrt{V_{\text{proxy}}}\sqrt{V_{\text{target}}}}}_{\rho(r^{\delta},r^{\Delta})}=\frac{1}{\sqrt{V_{\text{est}}}\sqrt{V_{\text{target}}}}=\text{IC}_{\text{final}}(\delta).$ (37)
@@ Appendix D Alternative Derivation via Partial Correlation Formula
In the main text, we derived the performance decomposition identity using a structural Linear Factor Model. In this appendix, we provide an alternative derivation based purely on the statistical properties of correlation. Specifically, we show that the decomposition can be rigorously understood as a special case of the Partial Correlation Formula where the residual term vanishes due to the specific signal dynamics of the market.
@@ D.1 The Decomposition and the Vanishing Residual
From a statistical perspective, the relationship between the model prediction $\hat{\mathbf{y}}_{t}^{\delta}$, the proxy label $\mathbf{r}_{t}^{\delta}$, and the final target $\mathbf{r}_{t}^{\Delta}$ can, under the standard linear/Gaussian framework, be expressed via the following correlation decomposition:
$\rho_{\hat{\mathbf{y}}_{t}^{\delta},\mathbf{r}_{t}^{\Delta}}=\underbrace{\rho_{\hat{\mathbf{y}}_{t}^{\delta},\mathbf{r}_{t}^{\delta}}\cdot\rho_{\mathbf{r}_{t}^{\delta},\mathbf{r}_{t}^{\Delta}}}_{\text{Mediated Path}}+\underbrace{\rho_{\hat{\mathbf{y}}_{t}^{\delta},\mathbf{r}_{t}^{\Delta}\cdot\mathbf{r}_{t}^{\delta}}\sqrt{1-\rho_{\hat{\mathbf{y}}_{t}^{\delta},\mathbf{r}_{t}^{\delta}}^{2}}\sqrt{1-\rho_{\mathbf{r}_{t}^{\delta},\mathbf{r}_{t}^{\Delta}}^{2}}}_{\text{Residual Term}}.$ (38)
@@
  • The Residual Term is controlled by the partial correlation $\rho_{\hat{\mathbf{y}}_{t},\mathbf{r}_{t}^{\Delta}\cdot\mathbf{r}_{t}^{\delta}}$. This term measures the correlation between the model’s prediction and the final target after the influence of the proxy $\mathbf{r}_{t}^{\delta}$ has been removed. Effectively, it asks: ”Can the model predict the return evolution from $\delta$ to $\Delta$ that is orthogonal to the return up to $\delta$?”
Application to Scenario 1 (Daily Close-to-Close): In Scenario 1, the input features are derived from history prior to the market open. Due to the high efficiency of the market, the predictive signal contained in these historical features is typically priced in almost immediately upon the open.
Consequently, the subsequent price movement from the intermediate horizon $\delta$ to the final target $\Delta$ is dominated by new, idiosyncratic information and noise that was not available at the decision time $t$. Since this future noise is strictly unforecastable based on the input features, the model’s predictive power for this residual component is negligible.
Mathematically, this implies $\rho_{\hat{\mathbf{y}}_{t},\mathbf{r}_{t}^{\Delta}\cdot\mathbf{r}_{t}^{\delta}}\approx 0$. Substituting this into Eq. (38), the entire residual term vanishes, and we recover the multiplicative decomposition presented in the main text:
$\rho_{\hat{\mathbf{y}}_{t}^{\delta},\mathbf{r}_{t}^{\Delta}}\approx\rho_{\hat{\mathbf{y}}_{t}^{\delta},\mathbf{r}_{t}^{\delta}}\cdot\rho_{\mathbf{r}_{t}^{\delta},\mathbf{r}_{t}^{\Delta}}.$ (39)

### v3

@@ 3.2 Theoretical Framework
The expected performance is determined by the net balance of two competing accumulation processes (ignoring a constant term):
@@
1. Information Gain reflects the growth of valid signal in the label. It increases as $\delta$ extends, but naturally saturates as $\alpha(\delta)\to const$ (when information is fully priced).
@@ 3.3 Generalization to Deep Learning
To empirically validate this theoretical connection, we specifically examine Scenario 1. Given the interday nature of this task, the market has ample time to absorb the prior day’s information, causing the signal embedded in the input features to be priced in immediately upon the market open. Consequently, the signal realization saturates almost immediately ($\alpha(\delta)\approx const$). This leads to the following approximation:
@@ 3.4 Mechanism Analysis
Scenario 1 (Interday): Interday signals typically saturate early (e.g., at the market open). Extending $\delta$ beyond this point yields vanishing information gain ($\alpha^{\prime}\approx 0$) while the noise penalty continues to grow linearly. Thus, the gradient is strictly negative, driving the optimal horizon to be much smaller than the target ($\delta^{*}\ll\Delta$).
@@ A.5.1 Predictive Model Training
A.5.2 Bi-Level Optimization Training.
@@ C.1.1 The Standard Static APT
$r^{\Delta}=\mathbf{w}^{*\top}\mathbf{s}+\epsilon^{\Delta},$ (16)
where $\mathbf{s}$ represents factor exposures, $\mathbf{w}^{*}$ represents factor risk premia, and $\epsilon^{\Delta}$ is the idiosyncratic noise. This model is typically static—it describes the return over a single, undefined period where information is assumed to be fully reflected.
@@ C.1.2 Temporal Extension
$r^{\delta}=\underbrace{\alpha(\delta)}_{\text{Signal Realization}}\mathbf{w}^{*\top}\mathbf{s}+\underbrace{\epsilon^{\delta}}_{\text{Accumulated Noise}}.$ (17)
@@
    Signal Realization Process ($\alpha(\delta)$): Unlike the standard APT, which assumes equilibrium (i.e., information is fully priced), we acknowledge that information incorporation takes time. Let $\alpha:(0,\Delta]\to(0,A]$ be a monotonically increasing function representing the degree of price discovery, where $A>0$ is a finite upper bound determined by the overall scale of the latent signal.
    • $\alpha(\delta)<A$ implies partial underreaction or gradual diffusion of information.
    • $\alpha(\delta)\approx A$ implies the signal is (effectively) fully realized at that horizon.
@@
    Here, $\sigma^{2}$ is the accumulation rate of idiosyncratic volatility, and $\delta_{0}>0$ represents intrinsic microstructure noise that exists even at $\delta\to 0$. While this assumption is generally reasonable at short horizons such as intraday (minute-level) and daily intervals, it may become less accurate over much longer horizons where structural breaks and slow-moving factors dominate. Since our study focuses on short-term stock forecasting at minute and daily frequencies, the random-walk-based noise model is well aligned with the forecasting regimes of interest.
@@ C.2 The Learning Setup: Finite-Sample OLS
$\hat{\mathbf{w}}_{\delta}=(\mathbf{S}^{\top}\mathbf{S})^{-1}\mathbf{S}^{\top}\mathbf{r}^{\delta}=\alpha(\delta)\mathbf{w}^{*}+(\mathbf{S}^{\top}\mathbf{S})^{-1}\mathbf{S}^{\top}\boldsymbol{\epsilon}^{\delta}.$ (19)
@@
$\hat{\mathbf{w}}_{\delta}=\alpha(\delta)\mathbf{w}^{*}+\mathbf{z}_{\delta},\quad\text{where }\mathbf{z}_{\delta}\sim\mathcal{N}\left(\mathbf{0},\frac{\sigma^{2}(\delta+\delta_{0})}{N}\mathbf{I}_{d}\right).$ (20)
@@ C.3 Derivation of Generalization Performance (Final IC)
    $r^{\Delta}=\alpha(\Delta)\mathbf{w}^{*\top}\mathbf{s}+\epsilon^{\Delta}.$
@@
    $\text{Cov}(\hat{y}^{\delta},r^{\Delta})$ $=\text{Cov}\big(\alpha(\delta)\mathbf{w}^{*\top}\mathbf{s}+\mathbf{z}_{\delta}^{\top}\mathbf{s},\,\alpha(\Delta)\mathbf{w}^{*\top}\mathbf{s}+\epsilon^{\Delta}\big)$ (21)
    $=\alpha(\delta)\alpha(\Delta)\,\text{Var}(\mathbf{w}^{*\top}\mathbf{s})+\underbrace{\text{Cov}(\mathbf{z}_{\delta}^{\top}\mathbf{s},\mathbf{w}^{*\top}\mathbf{s})}_{=0}+\underbrace{\text{Cov}(\alpha(\delta)\mathbf{w}^{*\top}\mathbf{s},\epsilon^{\Delta})}_{=0}+\underbrace{\text{Cov}(\mathbf{z}_{\delta}^{\top}\mathbf{s},\epsilon^{\Delta})}_{=0}$ (22)
    $=\alpha(\delta)\alpha(\Delta)\cdot 1=\alpha(\delta)\alpha(\Delta).$ (23)
@@
    $\text{Cov}(\hat{y}^{\delta},r^{\Delta})=\text{Cov}(\alpha(\delta)\mathbf{w}^{*\top}\mathbf{s},\alpha(\Delta)\mathbf{w}^{*\top}\mathbf{s})=\alpha(\delta)\alpha(\Delta).$ (24)
@@
    $V_{\text{target}}=\text{Var}(r^{\Delta})=\alpha(\Delta)^{2}+\sigma^{2}(\Delta+\delta_{0}).$ (26)
@@
$J(\delta)\triangleq\text{IC}_{\text{final}}^{2}(\delta)=\frac{\text{Cov}(\hat{y}^{\delta},r^{\Delta})^{2}}{V_{\text{est}}(\delta)V_{\text{target}}}=\frac{\alpha(\delta)^{2}\alpha(\Delta)^{2}}{\left[\alpha(\delta)^{2}+K(\delta+\delta_{0})\right]\cdot V_{\text{target}}},$ (27)
@@ C.4 Proof of the Paradox
To determine the optimal training horizon $\delta^{*}$, we analyze the behavior of the expected squared IC, $J(\delta)$. Since the target variance $V_{\text{target}}$ and $\alpha(\Delta)$ are constant scalars independent of the training horizon $\delta$, maximizing $J(\delta)$ is equivalent to maximizing the log-likelihood of $\alpha(\delta)^{2}/\left[\alpha(\delta)^{2}+K(\delta+\delta_{0})\right]$.
@@ C.5 Detailed Mechanism Analysis of Different Horizon Scenarios
Scenario 3: Intraday 90-minute Prediction (Hump-Shaped). When the target window extends to 90 minutes, we observe the characteristic Label Horizon Paradox. This scenario represents the transition between active information diffusion and signal saturation.
@@ C.6 Decomposition under Interday Prediction Scenario
Mathematically, this corresponds to the regime of rapid price discovery. The signal realization function satisfies $\alpha(\delta)\approx\text{const}$ and $\alpha^{\prime}(\delta)\approx 0$ for any horizon $\delta$ extending beyond the market open, and in particular $\alpha(\delta)\approx\alpha(\Delta)$ across the relevant horizons.
Under the general model $r^{\delta}=\alpha(\delta)\mathbf{w}^{*\top}\mathbf{s}+\epsilon^{\delta}$ and $r^{\Delta}=\alpha(\Delta)\mathbf{w}^{*\top}\mathbf{s}+\epsilon^{\Delta}$, and recalling that the idiosyncratic noise follows a nested random-walk structure such that $\text{Cov}(\epsilon^{\delta},\epsilon^{\Delta})=\text{Var}(\epsilon^{\delta})$, the covariance between the proxy label $r^{\delta}$ and the target $r^{\Delta}$ is
$\text{Cov}(r^{\delta},r^{\Delta})=\alpha(\delta)\alpha(\Delta)\,\text{Var}(\mathbf{w}^{*\top}\mathbf{s})+\text{Cov}(\epsilon^{\delta},\epsilon^{\Delta})=\alpha(\delta)\alpha(\Delta)+\sigma^{2}(\delta+\delta_{0}).$ (36)
The proxy and target variances are given by
$V_{\text{proxy}}(\delta)=\text{Var}(r^{\delta})=\alpha(\delta)^{2}+\sigma^{2}(\delta+\delta_{0}),$ (37)
$V_{\text{target}}=\text{Var}(r^{\Delta})=\alpha(\Delta)^{2}+\sigma^{2}(\Delta+\delta_{0}).$ (38)
The corresponding correlations are
$\rho(\hat{y}^{\delta},r^{\delta})=\frac{\text{Cov}(\hat{y}^{\delta},r^{\delta})}{\sqrt{V_{\text{est}}}\sqrt{V_{\text{proxy}}}}=\frac{\alpha(\delta)^{2}}{\sqrt{V_{\text{est}}}\sqrt{V_{\text{proxy}}}},$ (39)
$\rho(r^{\delta},r^{\Delta})=\frac{\text{Cov}(r^{\delta},r^{\Delta})}{\sqrt{V_{\text{proxy}}}\sqrt{V_{\text{target}}}}=\frac{\alpha(\delta)\alpha(\Delta)+\sigma^{2}(\delta+\delta_{0})}{\sqrt{V_{\text{proxy}}}\sqrt{V_{\text{target}}}},$ (40)
and
$\text{IC}_{\text{final}}(\delta)=\rho(\hat{y}^{\delta},r^{\Delta})=\frac{\text{Cov}(\hat{y}^{\delta},r^{\Delta})}{\sqrt{V_{\text{est}}}\sqrt{V_{\text{target}}}}=\frac{\alpha(\delta)\alpha(\Delta)}{\sqrt{V_{\text{est}}}\sqrt{V_{\text{target}}}}.$ (41)
Combining these, the final Information Coefficient admits the following general multiplicative form:
$\text{IC}_{\text{final}}(\delta)=\rho(\hat{y}^{\delta},r^{\delta})\,\rho(r^{\delta},r^{\Delta})\,\frac{\alpha(\delta)\alpha(\Delta)}{\alpha(\delta)^{2}}\cdot\frac{V_{\text{proxy}}(\delta)}{\alpha(\delta)\alpha(\Delta)+\sigma^{2}(\delta+\delta_{0})}.$ (42)
In the interday rapid price discovery regime, we have $\alpha(\delta)\approx\alpha(\Delta)$ for all relevant proxy horizons. Substituting this into the general expressions, the covariance between the proxy label and the target simplifies to
$\text{Cov}(r^{\delta},r^{\Delta})=\alpha(\delta)\alpha(\Delta)+\sigma^{2}(\delta+\delta_{0})\approx\alpha(\delta)^{2}+\sigma^{2}(\delta+\delta_{0})=V_{\text{proxy}}(\delta).$ (43)
Under this approximation, the proxy effectively behaves as a noisy but nearly unbiased version of the target at the same signal realization level, and the final IC is numerically close to the product of the model’s fit to the proxy and the proxy’s correlation with the target:
$\text{IC}_{\text{final}}(\delta)\;\approx\;\rho(\hat{y}^{\delta},r^{\delta})\times\rho(r^{\delta},r^{\Delta}).$ (44)
@@ Appendix D Alternative Derivation via Partial Correlation Formula
In the main text, we derived the performance decomposition identity using a structural Linear Factor Model. In this appendix, we provide an alternative derivation based purely on the statistical properties of correlation. Specifically, we show that the decomposition can be rigorously understood as a special case of the Partial Correlation Formula where the residual term vanishes due to the specific signal dynamics of the market. This shows that the multiplicative decomposition of the final IC is not an artifact of the specific factor-model setup, but a direct consequence of the vanishing partial correlation between the model prediction and the residual return beyond the proxy horizon. For notational brevity, we omit the decision time subscript $t$ and the stock index $i$ in the subsequent analysis.
@@ D.1 The Decomposition and the Vanishing Residual
From a statistical perspective, the relationship between the model prediction $\hat{y}^{\delta}$, the proxy label $r^{\delta}$, and the final target $r^{\Delta}$ can, under the standard linear/Gaussian framework, be expressed via the following correlation decomposition:
$\rho_{\hat{y}^{\delta},r^{\Delta}}=\underbrace{\rho_{\hat{y}^{\delta},r^{\delta}}\cdot\rho_{r^{\delta},r^{\Delta}}}_{\text{Mediated Path}}+\underbrace{\rho_{\hat{y}^{\delta},r^{\Delta}\cdot r^{\delta}}\sqrt{1-\rho_{\hat{y}^{\delta},r^{\delta}}^{2}}\sqrt{1-\rho_{r^{\delta},r^{\Delta}}^{2}}}_{\text{Residual Term}}.$ (45)
@@
  • The Residual Term is controlled by the partial correlation $\rho_{\hat{y}^{\delta},r^{\Delta}\cdot r^{\delta}}$. This term measures the correlation between the model’s prediction and the final target after the influence of the proxy $r^{\delta}$ has been removed. Effectively, it asks: Can the model predict the return evolution from $\delta$ to $\Delta$ that is orthogonal to the return up to $\delta$?
Application to Scenario 1 (Interday Prediction): In Scenario 1, the input features are derived from history prior to the market open. Due to the high efficiency of the market, the predictive signal contained in these historical features is typically priced in almost immediately upon the open.
Consequently, the subsequent price movement from the intermediate horizon $\delta$ to the final target $\Delta$ is dominated by new, idiosyncratic information and noise that was not available at the decision time. Since this future noise is strictly unforecastable based on the input features, the model’s predictive power for this residual component is negligible.
Mathematically, this implies $\rho_{\hat{y}^{\delta},r^{\Delta}\cdot r^{\delta}}\approx 0$. Substituting this into Eq. (45), the entire residual term vanishes, and we recover the multiplicative decomposition presented in the main text:
$\rho_{\hat{y}^{\delta},r^{\Delta}}\approx\rho_{\hat{y}^{\delta},r^{\delta}}\cdot\rho_{r^{\delta},r^{\Delta}}.$ (46)